# Harmonic covers of skeleta

### v1

@@ 8. Riemann-Hurwitz for skeleta
In the upcoming work of Hübner-Temkin [HT25] it will be shown that over any nonarchimedean field $k$ (not necessarily discretely valued or algebraically closed) the analytic different function $\delta$ associated to an étale morphism of $k$-analytic spaces defines a continuous function which is additionally PL on PL-charts when $k$ is stable. If we admit continuity of $\delta$, defined via the aforementioned formula, then by Lemma 8.6 and the fact that $\mathrm{sk}({X}^{\dagger})^{\mathrm{div}\,}$ is dense in $\mathrm{sk}({X}^{\dagger})$ (see [MN15, §2.4.12]) we see that the constructions agree: $\delta|_{\mathrm{sk}({X}^{\dagger})}=\delta_{X^{\prime\dagger}/{X}^{\dagger}}$
@@
By 7.32 we have $\mathrm{sp}_{*}\omega_{{X}^{\dagger}/S^{\dagger}}^{\otimes m}=mK_{\Sigma}$. Suppose ${X}$ is $\mathbb{Q}$-factorial, then by 6.20 it follows that if $\omega\in\Gamma({X}_{k},\omega_{{X}_{k}^{\otimes m}/k})\setminus\{0\}$ is a nonzero $m$-canonical form with $\mathrm{div}\,\omega$ supported on $D_{X}$ then
@@
By definition, the support function $F_{\mathrm{div}\,\omega}$ is equal to the weight function on $\Sigma$
$\mathrm{wt}_{\omega}:\Sigma\to\mathbb{R}.$
Now equip ${X}$ with the standard log-structure and let $\omega\in\Gamma({X}_{k},\omega_{{X}_{k}/k})\setminus\{0\}$. Then we obtain the formula
@@
$\rho_{*}\mathrm{div}\,\omega:=\sum_{(\sigma,\tau):\ \sigma\text{ unbounded},\ \tau\text{ bounded}}\partial_{\sigma/\tau}[\tau]$
summing over facet-ridge pairs with $\sigma$ unbounded and $\tau$ bounded. This generalises the results of [BN16, §3.2] to any dimension.
@@
Let $\omega\in\Gamma({X}_{k},\omega_{{X}_{k}^{\otimes m}/k})\setminus\{0\}$ be supported on $D_{{X}}$, then as functions on $\mathrm{sk}(X^{\prime\dagger})$ we have
@@ Appendix A Conventions on PL geometry
A morphisms of cones $\sigma\to\sigma^{\prime}$ is a group morphism $h:N_{\sigma}\to N_{\sigma^{\prime}}$ with $h_{\mathbb{R}}(|\sigma|)\subset|\sigma^{\prime}|$; note then $h_{\mathbb{R}}|_{|\sigma|}:|\sigma|\to|\sigma^{\prime}|$ is a morphism of $\mathbb{R}_{\geq 0}$-monoids. This defines the category of cones $\mathbf{Cone}$.
If $(\sigma,N_{\sigma})$ is a cone, then it is immediate that the monoid of integral points $\sigma\cap N_{\sigma}$ is integral, saturated and torsion-free. By Gordan’s lemma $\sigma\cap N_{\sigma}$ is also finitely generated101010proof: Write $N=N_{\sigma}$ and $\sigma=\sum_{i=1}^{k}\mathbb{R}_{\geq 0}m_{i}$ for some $m_{1},\dots,m_{k}\in\sigma$. Denote by $\|\cdot\|_{N}$ the Euclidean norm on $N_{\mathbb{R}}$ induced by a $\mathbb{Z}$-basis of $N$. Then every element in $\sigma\cap N$ is at most a distance $\epsilon=\sum_{i=1}^{k}\|m_{i}\|_{N}$ removed from an element in the submonoid $\sum_{i=1}^{k}\mathbb{N}m_{i}$ of $\sigma\cap N$. It follows that $\{m_{1},\dots,m_{k}\}$ and the collection of lattice points in the ball of radius $\epsilon$ around 00 jointly generate $M_{\sigma}$, as desired..
@@
We define a morphism of cone complexes $\phi:\Sigma\to\Sigma^{\prime}$ as a map $h:\tau\to\tau^{\prime}$ together with a family $(h_{\sigma}:\sigma\to h(\sigma))_{\sigma\in\tau}$ of compatible cone morphisms, where compatibility means that whenever $\rho$ is a face of $\sigma$, the restriction $(h_{\sigma})|_{\rho}$ is the composition of $h_{\rho}$ with the face embedding of $h(\rho)$ in $h(\sigma)$. Clearly $h$ induces a continuous map $|h|:|\Sigma|\to|\Sigma^{\prime}|$, which we will often identify with $h$ for simplicity.
@@
Arbitrary unions of ideals (resp. prime ideals) are again ideals (resp. prime ideals), and finite intersections of ideals are again ideals. If $\phi:M\to M^{\prime}$ is a morphism of monoids then the inverse image of a prime ideal is a prime ideal.
@@
Let $M$ be a monoid and let $S$ be a submonoid. Then the localisation $S^{-1}M$ is the quotient of $S\times M$ by the congruence relation generated by the relations $(s,m)\sim(s^{\prime},m^{\prime})$ for all $s,s^{\prime}\in S$ and $m,m^{\prime}\in M$ such that $sm^{\prime}=sm^{\prime}$. For $f\in M$ we write $M_{f}=\{f^{n}:n\in\mathbb{N}\}^{-1}M$. We have a canonical homeomorphism $D(f)=\mathrm{Spec}\,M_{f}$. For $\mathfrak{p}\in\mathrm{Spec}\,M$ we write $M_{\mathfrak{p}}=(M\setminus\mathfrak{p})^{-1}M$, and as a subset of $\mathrm{Spec}\,M$, we have that $\mathrm{Spec}\,M_{\mathfrak{p}})$ (resp. $\mathrm{Spec}\,(M/\mathfrak{p}\setminus\{0\})=\mathrm{Spec}\,M\setminus\mathfrak{p}$ (resp. are those primes $\mathfrak{q}$ contained in $\mathfrak{p}$ (resp. containing $\mathfrak{p}$).
@@
A $\mathbb{Z}$-PL polyhedral complex, or simply $\mathbb{Z}$-PL complex is a pair $\Sigma=(|\Sigma|,\tau)$ consisting of a locally compact topological space $|\Sigma|$ and a polyhedral atlas $\tau$, which is a set of polyhedrons such that (i)-(iv) holds:
@@
    For any two polyhedrons $\sigma_{1},\sigma_{2}\in\tau$, there is a unique polyhedron $\sigma_{1}\cap\sigma_{2}\in\tau$ such that $|\sigma_{1}\cap\sigma_{2}|=|\sigma_{1}|\cap|\sigma_{2}|$ and $\sigma_{1}\cap\sigma_{2}$ admits a face embedding in both $\sigma_{1}$ and $\sigma_{2}$.

### v2

@@ 8. Riemann-Hurwitz for skeleta
In the upcoming work of Hübner-Temkin [HT25] it will be shown that over any nonarchimedean field $k$ (not necessarily discretely valued or algebraically closed) the analytic different function $\delta$ associated to a generically étale morphism of $k$-analytic spaces defines a continuous function which is additionally PL on PL-charts when $k$ is stable. If we admit continuity of $\delta$, defined via the aforementioned formula, then by Lemma 8.6 and the fact that $\mathrm{sk}({X}^{\dagger})^{\mathrm{div}\,}$ is dense in $\mathrm{sk}({X}^{\dagger})$ (see [MN15, §2.4.12]) we see that the constructions agree: $\delta|_{\mathrm{sk}({X}^{\dagger})}=\delta_{X^{\prime\dagger}/{X}^{\dagger}}$
@@
By 7.32 we have $\mathrm{sp}_{*}\omega_{{X}^{\dagger}/S^{\dagger}}^{\otimes m}=mK_{\Sigma}$. Suppose ${X}$ is $\mathbb{Q}$-factorial, then by 6.20 it follows that if $\omega\in\Gamma({X}_{k},\omega_{{X}_{k}^{\otimes m}/k})\setminus\{0\}$ is a nonzero $m$-canonical form with $\mathrm{div}\,\omega$ supported on $D_{X}$. By the Poincaré-Lelong formula (Theorem 6.20) we have
@@
By definition, the support function $F_{\mathrm{div}\,\omega}$ is equal to the weight function
$\mathrm{wt}_{\omega}:\Sigma\to\mathbb{R}$
on $\Sigma$ Now equip ${X}$ with the standard log-structure and let $\omega\in\Gamma({X}_{k},\omega_{{X}_{k}/k})\setminus\{0\}$. Then we obtain the formula
@@
$\rho_{*}\mathrm{div}\,\omega:=\sum_{(\sigma,\tau):\ \sigma\text{ unbounded},\ \tau\text{ vertical}}\partial_{\sigma/\tau}[\tau]$
summing over facet-ridge pairs $(\sigma,\tau)$ with $\sigma$ unbounded and $\tau$ vertical. This generalises the results of [BN16, §3.2] to any dimension.
@@
Let $\omega\in\Gamma({X}_{k},\omega_{{X}_{k}/k}^{\otimes m})\setminus\{0\}$ be supported on $D_{{X}}$, then as functions on $\mathrm{sk}(X^{\prime\dagger})$ we have
@@ Appendix A Conventions on PL geometry
A morphism of cones $\sigma\to\sigma^{\prime}$ is a group morphism $h:N_{\sigma}\to N_{\sigma^{\prime}}$ with $h_{\mathbb{R}}(|\sigma|)\subset|\sigma^{\prime}|$; note then $h_{\mathbb{R}}|_{|\sigma|}:|\sigma|\to|\sigma^{\prime}|$ is a morphism of $\mathbb{R}_{\geq 0}$-monoids. This defines the category of cones $\mathbf{Cone}$.
If $(\sigma,N_{\sigma})$ is a cone, then it is immediate that the monoid of integral points $\sigma\cap N_{\sigma}$ is integral, saturated and torsion-free. By Gordan’s lemma $\sigma\cap N_{\sigma}$ is also finitely generated111111proof: Write $N=N_{\sigma}$ and $\sigma=\sum_{i=1}^{k}\mathbb{R}_{\geq 0}m_{i}$ for some $m_{1},\dots,m_{k}\in\sigma$. Denote by $\|\cdot\|_{N}$ the Euclidean norm on $N_{\mathbb{R}}$ induced by a $\mathbb{Z}$-basis of $N$. Then every element in $\sigma\cap N$ is at most a distance $\epsilon=\sum_{i=1}^{k}\|m_{i}\|_{N}$ removed from an element in the submonoid $\sum_{i=1}^{k}\mathbb{N}m_{i}$ of $\sigma\cap N$. It follows that $\{m_{1},\dots,m_{k}\}$ and the collection of lattice points in the ball of radius $\epsilon$ around 00 jointly generate $M_{\sigma}$, as desired..
@@
We define a morphism of cone complexes $\phi:\Sigma\to\Sigma^{\prime}$ as a map $h:\tau\to\tau^{\prime}$ together with a family $(h_{\sigma}:\sigma\to h(\sigma))_{\sigma\in\tau}$ of compatible cone morphisms, where compatibility means that whenever $\rho$ is a face of $\sigma$, the restriction $(h_{\sigma})|_{\rho}$ is the composition of $h_{\rho}$ with the face embedding of $h(\rho)$ in $h(\sigma)$. Then $h$ induces a continuous map $|h|:|\Sigma|\to|\Sigma^{\prime}|$ which we will often identify with $h$ for simplicity.
@@
Arbitrary unions of (prime) ideals are again (prime) ideals, and finite intersections of ideals are again ideals. If $\phi:M\to M^{\prime}$ is a morphism of monoids then the inverse image of a prime ideal is a prime ideal.
@@
Let $M$ be a monoid and let $S$ be a submonoid. Then the localisation $S^{-1}M$ is the quotient of $S\times M$ by the congruence relation generated by the relations $(s,m)\sim(s^{\prime},m^{\prime})$ for all $s,s^{\prime}\in S$ and $m,m^{\prime}\in M$ such that $sm^{\prime}=sm^{\prime}$. For $f\in M$ we write $M_{f}=\{f^{n}:n\in\mathbb{N}\}^{-1}M$. We have a canonical homeomorphism $D(f)=\mathrm{Spec}\,M_{f}$. For $\mathfrak{p}\in\mathrm{Spec}\,M$ we write $M_{\mathfrak{p}}=(M\setminus\mathfrak{p})^{-1}M$, and as subsets of $\mathrm{Spec}\,M$ the spectra $\mathrm{Spec}\,M_{\mathfrak{p}}$ and $\mathrm{Spec}\,(M/\mathfrak{p}\setminus\{0\})=\mathrm{Spec}\,M\setminus\mathfrak{p}$ are those primes $\mathfrak{q}$ contained in $\mathfrak{p}$ and containing $\mathfrak{p}$ respectively.
@@
A $\mathbb{Z}$-PL polyhedral complex, or simply $\mathbb{Z}$-PL complex is a pair $\Sigma=(|\Sigma|,\tau)$ consisting of a locally compact topological space $|\Sigma|$ and a polyhedral atlas $\tau$, which is a set of polyhedra such that (i)-(iv) holds:
@@
    For any two polyhedra $\sigma_{1},\sigma_{2}\in\tau$, there is a unique polyhedron $\sigma_{1}\cap\sigma_{2}\in\tau$ such that $|\sigma_{1}\cap\sigma_{2}|=|\sigma_{1}|\cap|\sigma_{2}|$ and $\sigma_{1}\cap\sigma_{2}$ admits a face embedding in both $\sigma_{1}$ and $\sigma_{2}$.